OpenStudy (anonymous):

How would i write this expression as the sine , cos, or tangent of a single angle? cos(π/5)cos(π/7)+sin(π/5)sin(π/7)

OpenStudy (anonymous):

\[\cos(A-B)\] Do you know the expansion of this?

OpenStudy (anonymous):

So it would be... Cos(π/5-π/7)?

OpenStudy (anonymous):

@koushik_ksv ?

OpenStudy (anonymous):

Yeah, even the other way round \[\cos (\Pi/7 - \Pi/5)\]. Its the same..

OpenStudy (anonymous):

ok so my final answer would be cos (pi/2)? Im not sure how to subtract radians

OpenStudy (anonymous):

what is \(\frac{1}{5}-\frac{1}{7}\) ?

OpenStudy (anonymous):

don't let the arithmetic fool you ignore the \(\pi\) and just subtract

OpenStudy (anonymous):

2/35.

OpenStudy (anonymous):

Oh so 2pi/35?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the final answer would be cos (2pi/5)...correct?

OpenStudy (anonymous):

if you meant "cos (2pi/35)" then yes

OpenStudy (anonymous):

oh yes right. But since the question asks of a single angle should i convert to degrees?

OpenStudy (anonymous):

@satellite73 , could you please help me really quick.

OpenStudy (anonymous):

forget degrees, they have nothing to do with this problem

OpenStudy (anonymous):

Ok thanks @satellite73

OpenStudy (anonymous):

\(\frac{2\pi}{35}\) is a number, but also a measure of an angle in radian